# Recursive variational Gaussian approximation with the Whittle likelihood for linear non-Gaussian state space models

### v1

@@ Abstract
Parameter inference for linear and non-Gaussian state space models is challenging because the likelihood function contains an intractable integral over the latent state variables. Exact inference using Markov chain Monte Carlo is computationally expensive, particularly for long time series data. Variational Bayes methods are useful when exact inference is infeasible. These methods approximate the posterior density of the parameters by a simple and tractable distribution found through optimisation. In this paper, we propose a novel sequential variational Bayes approach that makes use of the Whittle likelihood for computationally efficient parameter inference in this class of state space models. Our algorithm, which we call Recursive Variational Gaussian Approximation with the Whittle Likelihood (R-VGA-Whittle), updates the variational parameters by processing data in the frequency domain. At each iteration, R-VGA-Whittle requires the gradient and Hessian of the Whittle log-likelihood, which are available in closed form for a wide class of models. Through several examples using a linear Gaussian state space model and a univariate/bivariate non-Gaussian stochastic volatility model, we show that R-VGA-Whittle provides good approximations to posterior distributions of the parameters and is very computationally efficient when compared to asymptotically exact methods such as Hamiltonian Monte Carlo.
@@ 1 Introduction
State space models (SSMs) are a widely used class of hierarchical time series models employed in a diverse range of applications in fields such as epidemiology (e.g., Dukic et al.,, 2012), economics (e.g., Chan and Strachan,, 2023), finance (e.g., Kim et al.,, 1998), and engineering (e.g., Gordon et al.,, 1993). In an SSM, the observed data depend on unobserved, time-evolving latent state variables and unknown parameters. The observations are assumed to be independent given the latent states, the dynamics of the latent states are described through a Markov process, and often a simplifying assumption of linearity is made. Bayesian inference for parameters of linear non-Gaussian state space models is challenging since the likelihood function involves an intractable integral over the latent states.
A popular class of inference methods for non-Gaussian SSMs is Markov chain Monte Carlo (MCMC). Popular MCMC methods include particle MCMC (Andrieu et al.,, 2010) and Hamiltonian Monte Carlo (HMC, e.g., Neal,, 2011; Betancourt and Girolami,, 2015). While MCMC methods can provide samples from the exact posterior distribution, they tend to be computationally expensive.
@@
Variational Bayes (VB) methods are becoming increasingly popular for Bayesian inference for state space models (Zammit-Mangion et al.,, 2012; Gunawan et al.,, 2021; Frazier et al.,, 2023). VB approximates the posterior distribution of model parameters with a simpler and more tractable class of distributions constructed using so-called variational parameters, which are unknown and estimated via optimisation techniques. While VB methods tend to be more computationally efficient than MCMC methods, they provide approximate posterior distributions rather than samples from the exact posterior distributions. For a review of VB methods, see Blei et al., (2017).
@@
There are two main classes of VB methods: batch and sequential. A batch VB framework (Ong et al.,, 2018; Tan and Nott,, 2018; Gunawan et al.,, 2024) requires an optimisation algorithm to repeatedly process the entire dataset until variational parameters converge to a fixed point. A sequential VB framework, on the other hand, updates variational parameters using one observation at a time. Sequential methods only pass through the entire dataset once, and are thus ideal for settings with large datasets. Sequential VB methods are available for statistical models for independent data (Lambert et al.,, 2022), for classical autoregressive models for time-series data (Tomasetti et al.,, 2022), for generalised linear mixed models (Vu et al.,, 2024), and for SSMs where the state evolution model is linear and Gaussian (Beal,, 2003, Chapter 5; Zammit Mangion et al.,, 2011; Mangion et al.,, 2011). Campbell et al., (2021) proposed a sequential VB approach that is applicable to an SSM with non-linear state dynamics and non-Gaussian measurement errors, but this approach can be computationally expensive as it requires solving an optimisation problem and a regression task for each time step.
@@
Our article proposes a sequential VB approach to estimate parameters in linear non-Gaussian state space models. This approach builds on the Recursive Variational Gaussian Approximation for Latent variable models (R-VGAL) algorithm of Vu et al., (2024). We first transform time-domain data into frequency-domain data and then use the Whittle likelihood (Whittle,, 1953) in a sequential VB scheme. The logarithm of the Whittle likelihood function is a sum of the individual log-likelihood components in Fourier space, and is thus well-suited for use in sequential updating. We call our algorithm the Recursive Variational Gaussian Approximation with the Whittle likelihood (R-VGA-Whittle) algorithm. At each iteration, R-VGA-Whittle sequentially updates the variational parameters by processing one frequency or a block of frequencies at a time. Although sequential in nature, and requiring only one pass through the frequencies, we note that the R-VGA-Whittle algorithm requires the full dataset to be available in order to construct the periodogram, from which we obtain the frequency-domain data.
@@
We demonstrate the performance of R-VGA-Whittle in simulation studies involving a linear Gaussian state space model and univariate and bivariate non-Gaussian stochastic volatility models. We show that R-VGA-Whittle is accurate and is much faster than Hamiltonian Monte Carlo with the exact likelihood (HMC, Duane et al.,, 1987; Neal,, 2011; Hoffman and Gelman,, 2014) as well as HMC based on the Whittle likelihood; hereafter, we refer to these two methods as HMC-exact and HMC-Whittle, respectively. We subsequently apply R-VGA-Whittle to analyse exchange rate data.
@@ 2.1 State space models
In SSMs, the latent state variables $\{\mathbf{x}_{t}\}_{t=1}^{T}$ are characterised by an initial density ${\mathbf{X}_{1}\sim p(\mathbf{x}_{1}\mid{\bm{\theta}})}$ and state transition densities
@@ 2.2.1 Univariate Whittle likelihood
Here, the Whittle likelihood is evaluated by summing over the non-negative frequencies only, since $f(\omega_{k};\bm{\theta})$ and $I(\omega_{k})$ are both symmetric about the origin for real-valued time series. The term $\omega_{k}=0$ is not included since $I(0)=0$ when the time series is demeaned. The term $\omega_{k}=\pi$ is also omitted since it has a different distribution from the other frequencies and its influence is asymptotically negligible (Salomone et al.,, 2020; Villani et al.,, 2022).
@@ 2.3 R-VGA-Whittle
Vu et al., (2024) recently extended the Recursive Variational Gaussian Approximation (R-VGA) algorithm of Lambert et al., (2022) to estimate parameters in generalised linear mixed models (GLMMs). Their algorithm, called Recursive Variational Gaussian Approximation for Latent variable models (R-VGAL), is given in Algorithm 1, and described in detail in Section S1 of the online supplement. R-VGAL was successfully applied to grouped data where observations in the same group are correlated. The sequential algorithm requires the gradient and Hessian of the group-specific log-likelihood $\log p(\mathbf{y}_{i}\mid\bm{\theta})$ at each iteration ${i=1,\dots,N}$, where each $i$ corresponds to a group and $N$ is the total number of groups. In a GLMM setting where latent random effects from different groups are independent, these quantities can be unbiasedly estimated using Fisher’s and Louis’ identities (Cappé et al.,, 2005). However, the latent states in SSMs follow a Markov process and are hence temporally dependent; this dependence makes the computation of the gradient and Hessian of $\log p(\mathbf{y}_{t}\mid\bm{\theta})$, ${t=1,\dots,T}$, challenging. While these quantities can be estimated via particle filtering (e.g., Doucet and Shephard,, 2012), one would need to run a particle filter from the first time period to the current time period for every step in the sequential algorithm. This quickly becomes infeasible for long time series.
Algorithm 1 R-VGAL
Input: exchangeable observations $\mathbf{y}_{1},\dots,\mathbf{y}_{N}$, initial values $\bm{\mu}_{0}$ and $\mathbf{\Sigma}_{0}$.
@@
Set $q_{0}(\bm{\theta})=N(\bm{\mu}_{0},\mathbf{\Sigma}_{0})$.
for $i=1,\dots,N$ do
[…]
end for
[…]
@@
for $k=1,\dots,\left\lfloor\frac{T-1}{2}\right\rfloor$.
We also implement the damping approach of Vu et al., (2024) to mitigate instability that may arise when traversing the first few frequencies. Damping is done by splitting an update into $D$ steps, where $D$ is user-specific. In each step, the gradient and the Hessian of the log of the Whittle likelihood are multiplied by a “step size” $a=\frac{1}{D}$, and the variational parameters are updated $D$ times during one VB iteration. This has the effect of splitting a log-likelihood contribution into $D$ “parts”, and using them to update variational parameters one part at a time. We summarise R-VGA-Whittle in Algorithm 2.
@@
Set $q_{0}(\bm{\theta})=N(\bm{\mu}_{0},\mathbf{\Sigma}_{0})$.
@@
              Set $q_{k,l-1}(\bm{\theta})=N(\bm{\mu}_{k,l-1},\mathbf{\Sigma}_{k,l-1})$.
@@
         Set $\bm{\mu}_{k}=\bm{\mu}_{k,D},\mathbf{\Sigma}_{k}=\mathbf{\Sigma}_{k,D}$, $q_{k}(\bm{\theta})=N(\bm{\mu}_{k},\mathbf{\Sigma}_{k})$.
@@ 2.4 R-VGA-Whittle with block updates
Step 3: Set $q_{0}(\bm{\theta})=N(\bm{\mu}_{0},\mathbf{\Sigma}_{0})$.
@@ 2.5 HMC-Whittle
We now discuss the use of HMC with the Whittle likelihood for estimating parameters in SSMs. Suppose we want to sample from the posterior distribution $p(\bm{\theta}\mid\mathbf{y}_{1:T})$. We introduce an auxiliary momentum variable $\mathbf{r}$ of the same length as $\bm{\theta}$ and assume $\mathbf{r}$ comes from a density $N(\mathbf{0},\mathbf{M})$, where $\mathbf{M}$ is a mass matrix that is often a multiple of the identity matrix. The joint density of $\bm{\theta}$ and $\mathbf{r}$ is $p(\bm{\theta},\mathbf{r}\mid\mathbf{y})\propto\exp(-H(\bm{\theta},\mathbf{r}))$, where
@@
The gradient $\nabla_{\bm{\theta}}\log p(\mathbf{y}_{1:T}\mid\bm{\theta})$ on the right hand side of (16) is typically intractable in SSMs and needs to be estimated using a particle filter (e.g., Nemeth et al.,, 2016). However, if the exact likelihood is replaced with the Whittle likelihood, an approximation to ${\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})}$ can be obtained without the need for running an expensive particle filter repeatedly. That is, rather than targeting the true posterior distribution $p(\bm{\theta}\mid\mathbf{y}_{1:T})$, we target an approximation $\tilde{p}(\bm{\theta}\mid\mathbf{y}_{1:T})\propto p_{W}(\mathbf{y}_{1:T}\mid\bm{\theta})p(\bm{\theta})$, where $p_{W}(\mathbf{y}_{1:T}\mid\bm{\theta})\equiv\exp(l_{W}(\bm{\theta}))$ is the Whittle likelihood as defined in (6) for a univariate time series, or (10) for multivariate time series. Then $\mathcal{L}(\bm{\theta})$ in (14) is replaced by the quantity $\mathcal{L}_{W}(\bm{\theta})=\log\tilde{p}(\bm{\theta}\mid\mathbf{y}_{1:T})$, and the term $\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})$ in (15) is replaced with
@@
Algorithm 4 One step of the Leapfrog algorithm
@@
     Sample $\mathbf{r}^{(0)}\sim N(\mathbf{0},\mathbf{M})$.

### v2

@@ Abstract
Parameter inference for linear and non-Gaussian state space models is challenging because the likelihood function contains an intractable integral over the latent state variables. While Markov chain Monte Carlo (MCMC) methods provide exact samples from the posterior distribution as the number of samples go to infinity, they tend to have high computational cost, particularly for observations of a long time series. Variational Bayes (VB) methods are a useful alternative when inference with MCMC methods is computationally expensive. VB methods approximate the posterior density of the parameters by a simple and tractable distribution found through optimisation. In this paper, we propose a novel sequential variational Bayes approach that makes use of the Whittle likelihood for computationally efficient parameter inference in this class of state space models. Our algorithm, which we call Recursive Variational Gaussian Approximation with the Whittle Likelihood (R-VGA-Whittle), updates the variational parameters by processing data in the frequency domain. At each iteration, R-VGA-Whittle requires the gradient and Hessian of the Whittle log-likelihood, which are available in closed form for a wide class of models. Through several examples using a linear Gaussian state space model and a univariate/bivariate non-Gaussian stochastic volatility model, we show that R-VGA-Whittle provides good approximations to posterior distributions of the parameters and is very computationally efficient when compared to asymptotically exact methods such as Hamiltonian Monte Carlo.
@@ 1 Introduction
State space models (SSMs) are a widely used class of hierarchical time series models employed in a diverse range of applications in fields such as epidemiology (e.g., Dukic et al.,, 2012), economics (e.g., Chan and Strachan,, 2023), finance (e.g., Kim et al.,, 1998), and engineering (e.g., Gordon et al.,, 1993). In an SSM, the observed data depend on unobserved, time-evolving latent state variables and unknown parameters. The observations are assumed to be conditionally independent given the latent states, the dynamics of the latent states are described through a Markov process, and often a simplifying assumption of linearity of the state variables is made. When the state variables are non-Gaussian, Bayesian inference for the parameters of these SSMs is challenging since the likelihood function involves an intractable integral over the latent states.
A popular class of inference methods for non-Gaussian SSMs is Markov chain Monte Carlo (MCMC). While MCMC methods can provide samples from the posterior distribution that are asymptotically exact, they tend to be computationally expensive. An early application of MCMC to nonlinear, non-Gaussian state-space models (SSMs) is the work of Meyer et al., (2003), which combines the Laplace approximation (Laplace,, 1986) and the extended Kalman filter (Harvey,, 1990) to approximate the likelihood function. This approximation is then used within a Metropolis-Hastings algorithm to estimate parameters in a nonlinear, non-Gaussian stochastic volatility model. While this likelihood approximation is analytical, it is only exact for linear, Gaussian SSMs. An alternative approach is the pseudo-marginal Metropolis-Hastings algorithm developed by Andrieu and Roberts, (2009), which employs an unbiased likelihood estimate obtained through a particle filter. However, particle filters are computationally intensive and prone to particle degeneracy, especially when dealing with large datasets (see Kantas et al., (2009) for a detailed discussion of this issue). Both approaches typically require many iterations, each involving a full data pass with complexity $\textrm{O}(T)$, where $T$ is the number of observations, making them costly for long data sequences. Additionally, other alternatives, such as the Hamiltonian Monte Carlo (HMC) method  (Duane et al.,, 1987; Neal,, 2011; Hoffman and Gelman,, 2014), target the joint posterior distribution of states and parameters and typically require many iterations to achieve parameter convergence when $T$ is large and the state dimension is high.
@@
Variational Bayes (VB) methods are becoming increasingly popular for Bayesian inference for state space models (Zammit-Mangion et al.,, 2012; Gunawan et al.,, 2021; Frazier et al.,, 2023). VB approximates the posterior distribution of model parameters with a simpler and more tractable class of distributions constructed using so-called variational parameters, which are unknown and estimated via optimisation techniques. While VB methods tend to be more computationally efficient than MCMC methods, they provide approximate posterior distributions. For a review of VB methods, see Blei et al., (2017).
@@
There are two main classes of VB methods: batch and sequential. A batch VB framework (e.g., Tran et al.,, 2017; Tan and Nott,, 2018; Loaiza-Maya et al.,, 2022; Gunawan et al.,, 2024) requires an optimisation algorithm to repeatedly process the entire dataset until variational parameters converge to a fixed point. A sequential VB framework, on the other hand, updates variational parameters using one observation at a time. Sequential methods only pass through the entire dataset once, and are thus ideal for settings with large datasets. Sequential VB methods are available for statistical models for independent data (Lambert et al.,, 2022), for classical autoregressive models for time-series data (Tomasetti et al.,, 2022), for generalised linear mixed models (Vu et al.,, 2024), and for SSMs where the state evolution model is linear and Gaussian (Beal,, 2003, Chapter 5; Zammit Mangion et al.,, 2011; Mangion et al.,, 2011). Campbell et al., (2021) proposed a sequential VB approach that is applicable to an SSM with non-linear state dynamics and non-Gaussian measurement errors, but this approach can be computationally expensive as it requires solving an optimisation problem and a regression task for each time step.
@@
Our article proposes a sequential VB approach to estimate parameters in linear non-Gaussian state space models. This approach builds on the Recursive Variational Gaussian Approximation (R-VGA) algorithm of Lambert et al., (2022). We first transform time-domain data into frequency-domain data and then use the Whittle likelihood (Whittle,, 1953) in a sequential VB scheme. The logarithm of the Whittle likelihood function is a sum of the individual log-likelihood components in Fourier space, and is thus well-suited for use in sequential updating. We call our algorithm the Recursive Variational Gaussian Approximation with the Whittle likelihood (R-VGA-Whittle) algorithm. At each iteration, R-VGA-Whittle sequentially updates the variational parameters by processing one frequency or a block of frequencies at a time. Although sequential in nature, and requiring only one pass through the frequencies, we note that the R-VGA-Whittle algorithm requires the full dataset to be available in order to construct the periodogram, from which we obtain the frequency-domain data.
@@
We demonstrate the performance of R-VGA-Whittle in simulation studies involving a linear Gaussian state space model and univariate and bivariate non-Gaussian stochastic volatility models. We show that R-VGA-Whittle is accurate and is much faster than HMC with the exact likelihood as well as HMC based on the Whittle likelihood; hereafter, we refer to these two methods as HMC-exact and HMC-Whittle, respectively. We subsequently apply R-VGA-Whittle to analyse exchange rate data.
@@ 2.1 State space models
In SSMs, the $d$-dimensional latent state variables $\{\mathbf{x}_{t}\}_{t=1}^{T}$ are characterised by an initial density ${\mathbf{X}_{1}\sim p(\mathbf{x}_{1}\mid{\bm{\theta}})}$ and state transition densities
@@ 2.2.1 Univariate Whittle likelihood
Here, the Whittle likelihood is evaluated by summing over the non-negative frequencies only, since $f(\omega_{k};\bm{\theta})$ and $I(\omega_{k})$ are both symmetric about the origin for real-valued time series. The term $\omega_{k}=0$ is not included since $I(0)=0$ when the time series is de-meaned. The term $\omega_{k}=\pi$ is also omitted since it has a different distribution from the other frequencies and its influence is asymptotically negligible (Salomone et al.,, 2020; Villani et al.,, 2022).
@@ 2.3 R-VGA-Whittle
The original R-VGA algorithm of Lambert et al., (2022) is a sequential VB algorithm for parameter inference in statistical models where the observations are independent, such as linear or logistic regression models. Given independent and exchangeable observations $\{\mathbf{y}_{i}:i=1,\dots,N\}$, this algorithm sequentially approximates the “pseudo-posterior” distribution of the parameters given data up to the $i$th observation, $p(\bm{\theta}\mid\mathbf{y}_{1:i})$, with a Gaussian distribution $q_{i}(\bm{\theta})=\mathrm{Gau}(\bm{\mu}_{i},\mathbf{\Sigma}_{i})$. The updates for $\bm{\mu}_{i}$ and $\mathbf{\Sigma}_{i}$ are available in closed form, as seen in Algorithm 1. We describe this algorithm in more detail in Section S1 of the online supplement.
At each iteration ${i=1,\dots,N}$, the R-VGA algorithm requires the gradient and Hessian of the log-likelihood contribution from each observation, $\log p(\mathbf{y}_{i}\mid\bm{\theta})$. In a setting with independent observations, these likelihood contributions are easy to obtain; however, in SSMs where the latent states follow a Markov process and are hence temporally dependent, the log-likelihood function is intractable, as shown in Section 2.1. The gradient and Hessian of $\log p(\mathbf{y}_{t}\mid\bm{\theta})$, ${t=1,\dots,T}$, are hence also intractable. While these quantities can be estimated via particle filtering (e.g., Doucet and Shephard,, 2012), one would need to run a particle filter from the first time period to the current time period for every step in the sequential algorithm. This quickly becomes infeasible for long time series.
Algorithm 1 R-VGA (Lambert et al.,, 2022)
Input: independent observations $\mathbf{y}_{1},\dots,\mathbf{y}_{N}$, initial values $\bm{\mu}_{0}$ and $\mathbf{\Sigma}_{0}$.
Set $q_{0}(\bm{\theta})=\mathrm{Gau}(\bm{\mu}_{0},\mathbf{\Sigma}_{0})$.
for $i=1,\dots,N$ do
     $\bm{\mu}_{i}=\bm{\mu}_{i-1}+\mathbf{\Sigma}_{i}\mathbb{E}_{q_{i-1}}(\nabla_{\bm{\theta}}\log p(\mathbf{y}_{i}\mid\bm{\theta}))$
     $\mathbf{\Sigma}_{i}^{-1}=\mathbf{\Sigma}_{i-1}^{-1}-\mathbb{E}_{q_{i-1}}(\nabla_{\bm{\theta}}^{2}\log p(\mathbf{y}_{i}\mid\bm{\theta}))$
     Set $q_{i}(\bm{\theta})=\mathrm{Gau}(\bm{\mu}_{i},\mathbf{\Sigma}_{i})$.
end for
@@
We now extend the R-VGA algorithm of Lambert et al., (2022) to state space models by replacing the time domain likelihood with a frequency domain alternative: the Whittle likelihood. From (6) and (10) we see that the log of the Whittle likelihood is a sum over individual log-likelihood components, a consequence of the asymptotic independence of the periodogram frequency components (Shao and Wu, 2007a, ). We replace the quantities ${\nabla_{\bm{\theta}}\log p(\mathbf{y}_{i}\mid\bm{\theta})}$ and ${\nabla_{\bm{\theta}}^{2}\log p(\mathbf{y}_{i}\mid\bm{\theta})}$ in Algorithm 1 with the gradient and Hessian of the Whittle log-likelihood evaluated at each frequency. In the univariate case, the individual log-likelihood contribution at frequency $\omega_{k}$ is
@@
for $k=1,\dots,\left\lfloor\frac{T-1}{2}\right\rfloor$. We set $S$ following an experiment on the sensitivity of the algorithm to the choice of $S$; see Section S4 of the online supplement for more details.
We also implement the damping approach of Vu et al., (2024) to mitigate instability that may arise when traversing the first few frequencies. Damping is done by splitting an update into $D$ steps, where $D$ is user-specific. In each step, the gradient and the Hessian of the log of the Whittle likelihood are multiplied by a “step size” $a=\frac{1}{D}$, and the variational parameters are updated $D$ times during one VB iteration. This has the effect of splitting a log-likelihood contribution into $D$ “parts”, and using them to update variational parameters one part at a time. In practice, we set $D$ through experimentation, and choose it so that initial algorithmic instability is reduced with only a small increase in computational cost. We summarise R-VGA-Whittle in Algorithm 2.
@@
Set $q_{0}(\bm{\theta})=\mathrm{Gau}(\bm{\mu}_{0},\mathbf{\Sigma}_{0})$.
@@
              Set $q_{k,l-1}(\bm{\theta})=\mathrm{Gau}(\bm{\mu}_{k,l-1},\mathbf{\Sigma}_{k,l-1})$.
@@
         Set $\bm{\mu}_{k}=\bm{\mu}_{k,D},\mathbf{\Sigma}_{k}=\mathbf{\Sigma}_{k,D}$, $q_{k}(\bm{\theta})=\mathrm{Gau}(\bm{\mu}_{k},\mathbf{\Sigma}_{k})$.
@@ 2.4 R-VGA-Whittle with block updates
Step 3: Set $q_{0}(\bm{\theta})=\mathrm{Gau}(\bm{\mu}_{0},\mathbf{\Sigma}_{0})$.
@@ 2.5 HMC-Whittle
We now discuss the use of HMC with the Whittle likelihood for estimating parameters in SSMs. Suppose we want to sample from the posterior distribution $p(\bm{\theta}\mid\mathbf{y}_{1:T})$. We introduce an auxiliary momentum variable $\mathbf{r}$ of the same length as $\bm{\theta}$ and assume $\mathbf{r}$ comes from a density $\mathrm{Gau}(\mathbf{0},\mathbf{M})$, where $\mathbf{M}$ is a mass matrix that is often a multiple of the identity matrix. The joint density of $\bm{\theta}$ and $\mathbf{r}$ is $p(\bm{\theta},\mathbf{r}\mid\mathbf{y})\propto\exp(-H(\bm{\theta},\mathbf{r}))$, where
@@
The gradient $\nabla_{\bm{\theta}}\log p(\mathbf{y}_{1:T}\mid\bm{\theta})$ on the right hand side of (16) is typically intractable in SSMs and needs to be estimated using a particle filter (e.g., Nemeth et al.,, 2016). However, if the exact likelihood is replaced with the Whittle likelihood, an approximation to ${\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})}$ can be obtained without the need for running an expensive particle filter repeatedly. Rather than targeting the true posterior distribution $p(\bm{\theta}\mid\mathbf{y}_{1:T})$, HMC-Whittle proceeds by targeting an approximation $\tilde{p}(\bm{\theta}\mid\mathbf{y}_{1:T})\propto p_{W}(\mathbf{y}_{1:T}\mid\bm{\theta})p(\bm{\theta})$, where $p_{W}(\mathbf{y}_{1:T}\mid\bm{\theta})\equiv\exp(l_{W}(\bm{\theta}))$ is the Whittle likelihood as defined in (6) for a univariate time series, or (10) for multivariate time series. Then $\mathcal{L}(\bm{\theta})$ in (14) is replaced by the quantity $\mathcal{L}_{W}(\bm{\theta})=\log\tilde{p}(\bm{\theta}\mid\mathbf{y}_{1:T})$, and the term $\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})$ in (15) is replaced with
@@
Algorithm 4 One step of the Leapfrog algorithm, Leapfrog($\bm{\theta}^{*},\mathbf{r}^{*},\epsilon$)
@@
     Sample $\mathbf{r}^{(0)}\sim\mathrm{Gau}(\mathbf{0},\mathbf{M})$.